OpenStudy (anonymous):

Please help! For a statistics ANOVA question, I have found the F value, but am not sure how to find the P value?

OpenStudy (kirbykirby):

You can use an F-table and look for the F value (or approximate F-value) in the table according to the number of degrees of freedom in the numerator and denominator

OpenStudy (kirbykirby):

If you need a precise p-value then you can use software, or wolfram to get the exact p-value

OpenStudy (anonymous):

How do I find the degrees of freedom in the numerator and the denominator?

OpenStudy (anonymous):

I have 4 groups and a total of 20 observations

OpenStudy (kirbykirby):

is it just a one-way ANOVA?

OpenStudy (kirbykirby):

if it is, then the numerator degrees of freedom = 4-1 = 3 (i.e. # of groups - 1) and the denominator degrees of freedom = 20 - 4 = 16 (i.e. # of observations - # of groups)

OpenStudy (anonymous):

Okay thats what i thought..

OpenStudy (kirbykirby):

So once you have that, and know your specified level of significance (the \(\alpha\) level), then you can look it up in a table, and you should find the appropriate p-value (or range of p-values... you might find something like 0.05 < p < 0.1)

OpenStudy (anonymous):

thanks!